OpenStudy (brinazarski):

It has been estimated that 1000 curies of a radioactive substance introduced at a point on the surface of the open sea would spread over an area of 40,000 km2 in 40 days. Assuming that the area covered by the radioactive substance is a linear function of time t and is always circular in shape, express the radius r of the contamination in terms of t.

OpenStudy (brinazarski):

I don't even know where to start (except r(t)).... help?

OpenStudy (ookawaiioo):

oh, this can be tricky. let me work on it hold on

OpenStudy (ookawaiioo):

hey can I express (t) time in days?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

1000 meters per second is the answer

OpenStudy (ookawaiioo):

lol ill give you an equation and my work.

OpenStudy (ookawaiioo):

Radius = 25.23 Km (t) = days By day 40, you will have contaminated 40,000Km squared

OpenStudy (anonymous):

LOLOLOLOL this is taking forever.

OpenStudy (dumbcow):

Area is linear function...slope = 40,000/40 = 1000 \[A(t) = 1000t = \pi r(t)^{2}\] \[r(t) = \sqrt{\frac{1000t}{\pi}}\]